The intensity distribution of solar radiation peaks at a wavelength of approximately 500 nm. Assuming the Sun radiates like a black body, calculate the flux of photons leaving the sun in W/m^2

.
1
Expert's answer
2020-08-30T02:23:04-0400

Using Stefan Boltzmann law, find the flux of photons j: 

j = σT4, where constant σ  = 5.67*10-8 W/(m2*K4)

T is found from Wien displacement law: 

λmax = σw /T, where constant σw = 0.002898 m*K

λmax = 5*10-7 m = 0.002898/T, hence, T = 5796 K

The flux of photons is:

j = 5.67*10-8*(5796)4 = 6.4*107 W/m2